### v1

@@ Abstract
We developed a Python package GEHONG to mock the three-dimensional spectral data cube under the observation of an ideal telescope for the Integral Field Spectrograph of the Chinese Space Station Telescope (CSST-IFS). This package can generate one-dimensional spectra corresponding to local physical properties at specific positions according to a series of two-dimensional distributions of physical parameters of target sources. In this way, it can produce a spatially resolved spectral cube of the target source. Two-dimensional distributions of physical parameters, including surface brightness, stellar population, and line-of-sight velocity, can be modeled using the parametric model or based on real observational data and numerical simulation data. For the generation of one-dimensional spectra, we have considered four types of spectra, including the stellar continuum spectra, ionized gas emission lines, AGN spectra, and stellar spectra. That makes GEHONG able to mock various types of targets, including galaxies, AGNs, star clusters, and HII regions.
@@ 1 Introduction
To better understand the CSST-IFS capability, we have developed a Python package specifically to mock the three-dimensional (3D) spectroscopic data cubes of its main scientific targets. The package is named GEHONG (GEnerate tHe data Of iNtegral field spectrograph of Galaxy), which is dedicated to generating a wide variety of high spatial resolution ($0.1^{{}^{\prime\prime}}\times 0.1^{{}^{\prime\prime}}$) 3D data cubes of galactic objects while incorporating as many necessary physical processes as possible. The 3D data cube output of GEHONG mocks the ideal scientific data as observed by an IFS instrument on a perfect telescope, without including instrumental or observational effects such as seeing, cosmic rays, or background light. These instrumental and observational effects are incorporated by the another software of the CSST-IFS (Yan et al., in preparation), which uses the idealized data cubes generated by GEHONG as input to produce the corresponding raw CCD images of mock observations. These mock raw CCD images are then the key input of the scientific processing system of the CSST-IFS. Not only that, the 3D data cube output of GEHONG is also designed to be compatible with the input of the IFS exposure time calculator (ETC). With that, ETC can be used to quickly estimate the observation mode of the specific IFS targets and the corresponding signal-to-noise ratio, which is helpful for the scientific pre-research of the CSST-IFS. Obviously, GEHONG’s output can be applied not only to CSST-IFS-related research but also to other high spatial resolution IFS instruments. In fact, similar packages have been developed for specific IFS instruments, such as SIMSPIN(Harborne et al. 2020) and RealSim-IFS(Bottrell & Hani 2022) for SAMI, MaNGA, and MUSE. However, while these tools are tailored for these instruments, GEHONG remains a versatile tool that can be applied to simulate data for a wide range of IFS instruments, offering flexibility for research beyond the specific configurations of these surveys.
@@ 4 Two-Dimensional Maps
To model the surface brightness distribution of galaxies, GEHONG provides the map2d.sersic_map module, which generates a two-dimensional Sérsic profile (Sérsic 1963; Graham & Driver 2005). This parametric model can effectively describe a wide range of galaxy light distributions, from exponential disks to de Vaucouleurs bulges. An example of a Sérsic-based surface brightness map is shown in the left panel of Figure 5, with the corresponding input parameters listed in Table 3. A simple code example demonstrating the usage of this module is provided in Appendix A.2.
@@
For modeling the velocity field of rotating galaxies, GEHONG implements the map2d.tanh_map module. This module generates an axisymmetric rotation map based on a hyperbolic tangent rotation curve (Andersen & Bershady 2013), a common approximation for disk galaxies. The middle panel of Figure 5 shows an example of such a velocity map, with the input parameters summarized in Table 3. A code example for generating this type of velocity field is provided in Appendix A.2.
For galaxies with an inclination angle approximately given by $i\approx\arccos(1-q)$, where $q$ is the axis ratio, the line-of-sight velocity at a position $(x,y)$ on the galaxy plane is expressed as (van der Kruit & Allen 1978):
@@
where $V_{\text{max}}$ is the maximum rotational velocity, and $R_{t}$ is the turnover radius at which the rotation curve flattens (Andersen & Bershady 2013).
@@ Appendix A Usage of GEHONG
    stellar = spec1d.StellarContinuum(conf, stellar_tem, mag=17, sfh=2, ceh=-0.3,
    vel=30000, vdisp=150, ebv=0.1)
@@
    pl = spec1d.AGN_Powerlaw(conf, m5100=18, alpha=-1.5, vel=30000, ebv=0.1)
@@
    fe = spec1d.AGN_FeII(conf, hbeta_broad=800.0, r4570=0.4, ebv=0.1, vel=30000)
@@
    blr = spec1d.AGN_BLR(conf, hbeta_flux=800, hbeta_fwhm=5000.0, vel=30000, ebv=0.1)
@@
    nlr = spec1d.AGN_NLR(conf, nlr_temp, halpha=500, logz=-0.3, vel=30000,
    vdisp=400, ebv=0.1)
@@
    star = spec1d.SingleStar(conf, star_temp, mag=15, teff=8000, feh=-0.1,
    logg=3, vel=800, ebv=0.1)

### v2

@@ Abstract
We developed a Python package GEHONG to mock the three-dimensional spectral data cube under the observation of an ideal telescope for the Integral Field Spectrograph of the Chinese Space Station Survey Telescope (CSST-IFS). This package can generate one-dimensional spectra corresponding to local physical properties at specific positions according to a series of two-dimensional distributions of physical parameters of target sources. In this way, it can produce a spatially resolved spectral cube of the target source. Two-dimensional distributions of physical parameters, including surface brightness, stellar population, and line-of-sight velocity, can be modeled using the parametric model or based on real observational data and numerical simulation data. For the generation of one-dimensional spectra, we have considered four types of spectra, including the stellar continuum spectra, ionized gas emission lines, AGN spectra, and stellar spectra. That makes GEHONG able to mock various types of targets, including galaxies, AGNs, star clusters, and HII regions.
@@ 1 Introduction
To better understand the CSST-IFS capability, we have developed a Python package specifically to mock the three-dimensional (3D) spectroscopic data cubes of its main scientific targets. The package is named GEHONG (GEnerate tHe data Of iNtegral field spectrograph of Galaxy), which is dedicated to generating a wide variety of high spatial resolution ($0.1^{{}^{\prime\prime}}\times 0.1^{{}^{\prime\prime}}$) 3D data cubes of galactic objects while incorporating as many necessary physical processes as possible. The 3D data cube output of GEHONG mocks the ideal scientific data as observed by an IFS instrument on a perfect telescope, without including instrumental or observational effects such as seeing, cosmic rays, or background light. These instrumental and observational effects are incorporated by the another software of the CSST-IFS (Yan et al. 2026), which uses the idealized data cubes generated by GEHONG as input to produce the corresponding raw CCD images of mock observations. These mock raw CCD images are then the key input of the scientific processing system of the CSST-IFS. Not only that, the 3D data cube output of GEHONG is also designed to be compatible with the input of the IFS exposure time calculator (ETC). With that, ETC can be used to quickly estimate the observation mode of the specific IFS targets and the corresponding signal-to-noise ratio, which is helpful for the scientific pre-research of the CSST-IFS. Obviously, GEHONG’s output can be applied not only to CSST-IFS-related research but also to other high spatial resolution IFS instruments. In fact, similar packages have been developed for specific IFS instruments, such as SIMSPIN(Harborne et al. 2020) and RealSim-IFS(Bottrell and Hani 2022) for SAMI, MaNGA, and MUSE. However, while these tools are tailored for these instruments, GEHONG remains a versatile tool that can be applied to simulate data for a wide range of IFS instruments, offering flexibility for research beyond the specific configurations of these surveys.
@@ 4 Two-Dimensional Maps
To model the surface brightness distribution of galaxies, GEHONG provides the map2d.sersic_map module, which generates a two-dimensional Sérsic profile (Sérsic 1963; Graham and Driver 2005). This parametric model can effectively describe a wide range of galaxy light distributions, from exponential disks to de Vaucouleurs bulges. An example of a Sérsic-based surface brightness map is shown in the left panel of Figure 5, with the corresponding input parameters listed in Table 3. A simple code example demonstrating the usage of this module is provided in Appendix A.2.
@@
For modeling the velocity field of rotating galaxies, GEHONG implements the map2d.tanh_map module. This module generates an axisymmetric rotation map based on a hyperbolic tangent rotation curve (Andersen and Bershady 2013), a common approximation for disk galaxies. The middle panel of Figure 5 shows an example of such a velocity map, with the input parameters summarized in Table 3. A code example for generating this type of velocity field is provided in Appendix A.2.
For galaxies with an inclination angle approximately given by $i\approx\arccos(1-q)$, where $q$ is the axis ratio, the line-of-sight velocity at a position $(x,y)$ on the galaxy plane is expressed as (van der Kruit and Allen 1978):
@@
where $V_{\text{max}}$ is the maximum rotational velocity, and $R_{t}$ is the turnover radius at which the rotation curve flattens (Andersen and Bershady 2013).
@@ Appendix A Usage of GEHONG
    stellar = spec1d.StellarContinuum(conf, stellar_tem, mag=17, sfh=2,
    ceh=-0.3, vel=30000, vdisp=150, ebv=0.1)
@@
    pl = spec1d.AGN_Powerlaw(conf, m5100=18, alpha=-1.5, vel=30000,
    ebv=0.1)
@@
    fe = spec1d.AGN_FeII(conf, hbeta_broad=800.0, r4570=0.4,
    ebv=0.1, vel=30000)
@@
    blr = spec1d.AGN_BLR(conf, hbeta_flux=800, hbeta_fwhm=5000.0,
    vel=30000, ebv=0.1)
@@
    nlr = spec1d.AGN_NLR(conf, nlr_temp, halpha=500, logz=-0.3,
    vel=30000, vdisp=400, ebv=0.1)
@@
    star = spec1d.SingleStar(conf, star_temp, mag=15, teff=8000,
    feh=-0.1, logg=3, vel=800, ebv=0.1)